# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v3

@@ Appendix D Direct Decision-based Attack Scheme
We start with describing our attack scheme as follows.
Step 1: Initial detection and BH labeling. Using the training data $\{Z_{j}\}_{j=1}^{n}$ and the mixed sample $\{Z_{j}\}_{j=k+1}^{n+m}$, form the dataset
$\mathcal{D}=\{(Z_{i},Y_{i})\}_{i=1}^{n+m}$
where $Y_{i}=0$ for $i\in[1:k]$ and $Y_{i}=1$ for $i\in[k+1:n+m]$ using the positive-unlabeled (PU) framework. Train the score function
$s(z)\leftarrow\mathrm{TrainScoreFunction}(\mathcal{D}).$
Note that $s(z)$ automatically satisfies the condition in equation II-A as $Y_{i}$ for $i\in[k+1:n+m]$ are the same. Compute empirical $p$-values for $i\in[1:m]$,
@@
Step 2: Attack set selection. Within the set of test samples (i.e., with indices from $[n+1:n+m]$), select a subset $\{Z_{n+i}:i\in\mathcal{A}\}$ from the unrejected test samples as the attack target.
We set the attack size as (1) fixed size where $|\mathcal{A}|=m_{a}$ for some fixed number $m_{a}$, or (2) random size where $m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$, where $\gamma\in(0,1]$ is an “attack intensity” parameter specified by the attacker. One natural choice of $\mathcal{A}$ is to select the unrejected hypotheses with the smallest $p$-values (i.e., those closest to the rejection boundary). Let $(i_{1},i_{2},\ldots,i_{m-R})$ denote the indices of unrejected hypotheses ordered by their $p$-values: $\hat{p}_{i_{1}}\leq\hat{p}_{i_{2}}\leq\cdots\leq\hat{p}_{i_{m-R}}$. Then we define $\mathcal{A}=\{i_{1},i_{2},\ldots,i_{m_{\mathcal{A}}}\}$. This selects the $m_{\mathcal{A}}$ unrejected indices with the smallest $p$-values, targeting hypotheses that are close to $\hat{\tau}$ and making it an effective attack strategy as demonstrated in our experiments.
@@
$\widetilde{Z}_{n+i}$ $=f_{\text{attack}}(Z_{n+i};\,s(z))$ (24)
$:=f_{\text{attack}}(Z_{n+i};\,\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\},(Z_{n+j}:j\in A\cup\mathcal{H}_{1}),(Z_{1},\dots,Z_{k}))$ (25)
such that $\mathbf{1}\{s(Z_{n+i})\geq 0.5\}\neq\mathbf{1}\{s(\widetilde{Z}_{n+i})\geq 0.5\}$, meaning that the decision is altered. Here we write
$\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$
as an unordered set to highlight that $f_{\text{attack}}$ does not depend on the order of elements in this set.
Step 4: Applying AdaDetect on the attacked data. After the attack, the user applies AdaDetect and computes the score function as the first step. As the data is now changed by the attacker, we denote the score function after the attack by $\tilde{s}$, and the empirical $p$-value after the attack by ${\tilde{p}}_{i}$ for $i\in[1:m]$. We stress that $\tilde{s}$ still satisfies equation II-A.
@@
$f_{\text{attack}}$ does not depend on the order of elements in $\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$.
This proposition captures the main subtle yet important difference between this attack scheme and the oracle setting in Theorem 1. It holds because $f_{\text{attack}}(\cdot\,;\,s(z))$ only relies on the score function $s(z)$, and $s(z)$ is invariant to order of elements in $\{Z_{k+1},\dots,Z_{n+m}\}$ according to equation II-A, as a consequence of the PU framework.
D-A Analysis
In this setting, we denote the corresponding FDR as $\text{FDR}^{*}_{\text{attack--decision}}$, and quantify the loss in FDR caused by the attack.
Theorem 2.
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation II-A and the attack scheme $f_{attack}$ being order-invariant as in equation 25, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (26)
@@ D-A Analysis
It has been proved that AdaDetect has a strong detection power (the probability of correctly rejecting a non-null), as shown in Theorem 5.1 from [7]. This implies that the set $\mathcal{A}$ will nearly contain all indices from true nulls because the non-nulls are mostly rejected. We use this to show that the upper bound $\mathbb{E}[\frac{m_{\mathcal{A}}}{{\widetilde{R}}_{m_{a}}\lor 1}]$ in equation 1 is relatively tight (see details in Proposition 3).
Proof of Theorem 2.
It follows directly from proof of Theorem 1, as the only difference between Theorem 2 and Theorem 1 is that the score function is trained differently. But according to Proposition 2, the score function for attack is still invariant under the permutation of $\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$, which makes the rest of the proof exactly the same as that for Theorem 1. ∎
@@
Consider that $\mathcal{A}$ is random with fixed size $|\mathcal{A}|=m_{a}$. Under Assumption 1, with equation II-A and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and ${\widetilde{R}}_{m_{a}}$.
Unlike the original AdaDetect, which guarantees $\text{FDR}\leq\alpha$ in benign settings, Corollary 1 provides an FDR upper-bound under adversarial perturbations. We also have the following corollary about the setting when $\mathcal{A}$ is random with a fixed size $m_{\mathcal{A}}$.
Corollary 2.
Consider that $\mathcal{A}$ is random with a random size $|\mathcal{A}|=m_{\mathcal{A}}$. Under Assumption 1, with equation II-A and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\alpha+\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (27)
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and ${\widetilde{R}}_{m_{a}}$.
In the following proposition, we show that the upper bound in equation 20 is relatively tight by making a connection between step (a) in equation 20 and the power of AdaDetect. Roughly speaking, the upper bound is relatively tight when the power of AdaDetect is decent. Instead of showing detailed technical steps following the proof of Theorem 5.1 from [7], we choose to provide a high-level argument to connect the power of AdaDetect and our upper bound in equation 20.
@@
Under the assumptions in Theorem 5.1 from [7] and when $\mathcal{A}$ is randomly selected from the unrejected indices with $m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$, we have that for some small $\delta^{\prime}$ and $\eta^{\prime}$,
@@
According to Theorem 5.1 from [7], we have that the rejection set by AdaDetect at level $\lambda$, denoted by $\text{AdaDetect}_{\lambda}$ satisfies
@@
Appendix E Comparison between our surrogate algorithm and INCREASE-c [18]
@@
To evaluate the comparative impact, we replicate the INCREASE-c experimental framework on the same credit card dataset. An isolation forest is trained with a training sample selected uniformly at random from the set of true nulls; a calibration subset of strictly genuine transactions is randomly selected from the null data, after which a test sample is formed by combining the remaining null data with a random subset of fraudulent transactions. The test sample is transformed to $p$-values, and the BH procedure (with a level of $0.1$) is applied to identify the set of fraudulent transactions. (The full details of the experiment can be found in [18].) We execute INCREASE-$c$ against these $p$-values and their ground-truth labels across $100$ simulations for each attack budget $c=1,5,10,20$. The results are reported in Table $6$. The high average value of the selected $p$-values (i.e., $\mathbb{E}[p_{\text{selected}}]$) confirms that INCREASE-c targets data points far from the (average) BH cutoff (i.e., the last column in Table 6). Due to the drastic change in perturbed $p$-values, there is a clear impact on the resulting FDR and total rejection count.
TABLE VI: Experiment E.1: INCREASE-c results
$c$ original FDR INCREASE-c FDR original $\mathbb{E}[R]$ INCREASE-c $\mathbb{E}[R]$ $\mathbb{E}[p_{\text{selected}}]$ avg. BH cutoff
@@
Appendix F Additional Experiments
Synthetic data generation.
We generate two types of data: null samples from distribution $P_{0}$ and non-null samples from distribution $P_{1}$. We let $d=20$ for all the synthetic data.
Independent Gaussian: We consider
$P_{0}=\mathcal{N}(0,I_{d}),\quad P_{1}=\mathcal{N}(\mu,I_{d}),$
where $\mu\in\mathbb{R}^{d}$ is a sparse mean shift vector: the first five coordinates are set to $\sqrt{2\log(d)}$ and the remaining coordinates are zero.
Non-Gaussian: We let the first two coordinates of nulls and non-nulls be drawn independently from Beta distributions, with
$P_{0}:(X_{1},X_{2})\sim\text{Beta}(5,5),$
$P_{1}:(X_{1},X_{2})\sim\text{Beta}(1,3).$
The remaining coordinates are drawn i.i.d. from $\text{Beta}(1,1)$ under both $P_{0}$ and $P_{1}$.
@@ Appendix F Additional Experiments
$\Sigma=c\mathbf{1}\mathbf{1}^{\top}+(b^{2}-c)I_{d},$
where $\mathbf{1}=[1,\dots,1]^{\top}\in\mathbb{R}^{d}$.
We define the null and non-null distributions as
$P_{0}=\mathcal{N}(a\mathbf{1},\Sigma),\qquad P_{1}=\mathcal{N}((a+\delta)\mathbf{1},\Sigma),$
where $\delta>0$ introduces a mean shift across all coordinates. Thus $P_{0}$ and $P_{1}$ share the same exchangeable covariance structure but differ in their mean vectors.
We consider the following four real-world datasets.
@@
    Credit Card: Transactions made by credit cards over two days, some of which are fraudulent [56].
  • $\bullet$
    KDDCup99: A set of network connections that includes a variety of simulated intrusions [57].
  • $\bullet$
    Mammography: Features extracted from mammograms, some with microcalcifications [58].
Experiment A.1: Varying attack size with random forest (RF) models.
We evaluate the impact of attack size on both oracle and surrogate attack performance using identical RF architectures for both score functions. The attack size $m_{a}$ sweeps between $50$ and $200$. We assess how each attack scheme performs under different attack scales with homogeneous RF configurations.
We only tested HSJA in synthetic data experiments while both of HSJA and Boundary attack will be evaluated in real-world data experiments. The result shows that the FDR for both attack schemes are under the theoretical upper bound, and oracle attack outperforms surrogate attack in non-gaussian case when the original power is low. Overall, both attack schemes successfully increase the FDR across varying attack sizes.
TABLE VII: Experiment A.1: FDR + RF
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original FDR $08\pm 0.03 0.0$$8\pm 0.04 0.0$$08\pm 0.04$
oracle ($m_{a}=50$) $0.36\pm 0.02$ $0.40\pm 0.05$ $0.38\pm 0.02$
surrogate ($m_{a}=50$) $0.34\pm 0.02$ $0.20\pm 0.05$ $0.37\pm 0.02$
estimated upper bound $0.43$ $0.42$ $0.41$
oracle ($m_{a}=200$) $0.67\pm 0.00$ $0.71\pm 0.01$ $0.69\pm 0.00$
surrogate ($m_{a}=200$) $0.67\pm 0.00$ $0.64\pm 0.01$ $0.67\pm 0.00$
estimated upper bound $0.80$ $0.78$ $0.77$
TABLE VIII: Experiment A.1: Power + RF
Dataset Independent Gaussian Non-Gaussian Exchangeable Gaussian
original power $0.96\pm 0.02$ $0.55\pm 0.06$ $1.00\pm 0.00$
oracle ($m_{a}=50$) $0.99\pm 0.01$ $0.87\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=50$) $0.96\pm 0.02$ $0.65\pm 0.05$ $1.00\pm 0.00$
oracle ($m_{a}=200$) $0.99\pm 0.01$ $0.98\pm 0.01$ $1.00\pm 0.00$
surrogate ($m_{a}=200$) $0.96\pm 0.02$ $0.78\pm 0.05$ $1.00\pm 0.00$
Experiment A.2: Mismatched score function configurations.
We investigate both oracle and surrogate attack performance when using different model architectures and parameters for the score functions, with attack size $m_{a}=200$. This experiment comprises distinct configurations for each attack type:
  • $\bullet$
    RF–NN: AdaDetect score function $s(z)$ uses an RF, attack model score function g(z){g}(z) uses an NN.
  • $\bullet$
    RF–RF: Both score functions use RFs with same hyperparameters.
TABLE IX: Experiment A.2: FDR + RF-NN
@@
TABLE XI: Experiment A.3: Power + NN
@@
Experiment A.3: Real-world data with NN models.
This experiment employs NN architectures for both score functions on the four real-world datasets, evaluating both oracle and surrogate attack performance with attack size $m_{a}=200$. This allows us to evaluate how each attack type performs when both the target model and attacker use NN-based approaches on realistic data distributions. We compare boundary attack with our default HSJA attack under both oracle and surrogate attack schemes, using identical RF architectures for both score functions with $m_{a}=200$. The result shows that both HSJA and boundary attack are successful at rasing the FDR for NN models in real-world data.

### v4

@@ Appendix B Direct Decision-based Attack Scheme
Step 1: Initial detection and BH labeling. Using the training data $\{Z_{j}\}_{j=1}^{n}$ and the mixed sample $\{Z_{j}\}_{j=k+1}^{n+m}$, form the dataset $\mathcal{D}=\{(Z_{i},Y_{i})\}_{i=1}^{n+m}$, where $Y_{i}=0$ for $i\in[1:k]$ and $Y_{i}=1$ for $i\in[k+1:n+m]$ using the positive-unlabeled (PU) framework. Train the score function $s(z)\leftarrow\mathrm{TrainScoreFunction}(\mathcal{D})$. Note that $s(z)$ automatically satisfies the condition in equation 3 as $Y_{i}$ for $i\in[k+1:n+m]$ are the same. Compute empirical $p$-values for $i\in[1:m]$,
@@
Step 2: Attack set selection. Within the set of test samples (i.e., with indices from $[n+1:n+m]$), select a subset $\{Z_{n+i}:i\in\mathcal{A}\}$ from the unrejected test samples as the attack target. We set the attack size as (1) fixed size where $|\mathcal{A}|=m_{a}$ for some fixed number $m_{a}$, or (2) random size with size $m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$, where $\gamma\in(0,1]$ is an “attack intensity” parameter specified by the attacker. However, we have not implemented this random set setting in our experiments.
@@
$\widetilde{Z}_{n+i}$ $=f_{\text{attack-decision}}(Z_{n+i};\,s(z))$ (27)
$:=f_{\text{attack-decision}}(Z_{n+i};\,\{Z_{1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\},(Z_{n+j}:j\in A\cup\mathcal{H}_{1}))$ (28)
such that $\mathbf{1}\{s(Z_{n+i})\geq 0.5\}\neq\mathbf{1}\{s(\widetilde{Z}_{n+i})\geq 0.5\}$, meaning that the decision is altered.
Step 4: Applying AdaDetect on the attacked data. After the attack, the user applies AdaDetect and computes the score function as the first step. As the data is now changed by the attacker, we denote the score function after the attack by $\tilde{s}(z)$, and the empirical $p$-value after the attack by ${\tilde{p}}_{i}$ for $i\in[1:m]$.
@@
$f_{\text{attack-decision}}(\cdot\,;\,g(z))$ does not depend on the order of elements in $\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$.
This proposition captures the main subtlety difference between this attack scheme and the oracle setting in Theorem 1. It holds because $f_{\text{attack-decision}}(\cdot\,;\,g(z))$ only relies on the score function $s(z)$, and $s(z)$ is invariant to order of elements in $\{Z_{k+1},\dots,Z_{n+m}\}$ according to equation 3, as a consequence of the PU framework.
Proposition 3.
Consider that $\mathcal{A}$ is a fixed set with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation 3 and the attack scheme $f_{\text{attack-decision}}(\cdot\,;\,g(z))$ being order-invariant as in equation 28, the FDR after the attack, denoted by $\text{FDR}^{*}_{\text{attack--decision}}$ can be upper bounded as
$\text{FDR}^{*}_{\text{attack--decision}}\leq\frac{m_{0}}{m}\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (29)
@@
When $\mathcal{A}$ is random with fixed size $|\mathcal{A}|=m_{a}$, we have $\text{FDR}_{\text{attack}}\leq\text{FDR}^{*}_{\text{attack--decision}}$.
Furthermore, we have $\text{FDR}_{\text{attack}}\leq\text{FDR}^{*}_{\text{attack--decision}}\leq\alpha$ as long as
$h(m_{a})\leq\frac{m_{1}}{m}\,\alpha\,\quad\text{ where }h(x):=m_{a}\,\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}(m_{a})\lor 1}\right].$ (30)
@@
The reason we have $m_{0}/m$ in equation 29 rather than $(m_{0}-m_{a})/m$ as in Theorem 1 is because $|\mathcal{H}_{0}\setminus\mathcal{A}|$ may be strictly bigger than $m_{0}-m_{a}$ but still bounded by $m_{0}$, while it is exactly equal to $m_{0}-m_{a}$ in the oracle setting.
This result follows directly from proof of Theorem 1, as the only difference between Theorem 3 and Proposition 1 is that the score function is trained differently. But according to Proposition 2, the score function for attack is still invariant under the permutation of $\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$, which implies that the “matching” of the invariance between $f_{\text{attack-decision}}(\cdot\,;\,g(z))$ and $\tilde{s}(z)$ still holds. Another way to see this is that because both the attacker and user apply PU, and thus assign the same labels to $\{Z_{k+1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$. This makes the rest of the proof exactly the same as that for Theorem 1.
When the size of the attack set is random, denoted by $m_{\mathcal{A}}$, it will show up inside the expectation as follows.
@@
Consider that $\mathcal{A}$ is random with a random size $|\mathcal{A}|=m_{\mathcal{A}}$. Under Assumption 1, with equation 3 and equation 7, the FDR after the attack is
$\text{FDR}_{\text{attack}}\leq\frac{m_{0}}{m}\alpha+\mathbb{E}\left[\frac{m_{\mathcal{A}}}{{\widetilde{R}}(m_{\mathcal{A}})\lor 1}\right],$ (31)
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$, which induces randomness in $\mathcal{A}$ and ${\widetilde{R}}(m_{\mathcal{A}})$.
B-A Attack set selection
It has been proved that AdaDetect has a strong detection power (the probability of correctly rejecting a non-null), as shown in [7, Theorem 5.1]. This implies that, in the surrogate method, the set $\mathcal{A}$ will nearly contain all indices from true nulls because the non-nulls are mostly rejected. We use this to show that $\mathrm{P}(\mathcal{A}\,\subseteq\,\mathcal{H}_{0})$ is high. Instead of showing detailed technical steps following the proof [7, Theorem 5.1], we choose to provide an informal argument to connect the power of AdaDetect and $\mathrm{P}(\mathcal{A}\,\subseteq\,\mathcal{H}_{0})$.
@@ B-A Attack set selection
Under the assumptions in [7, Theorem 5.1] and when $\mathcal{A}$ is randomly selected from the unrejected indices with $m_{\mathcal{A}}$, we have that for some small $\delta^{\prime}$ and $\eta^{\prime}$,
@@
According to from [7, Theorem 5.1], we have that the rejection set by AdaDetect at level $\lambda$, denoted by $\text{AdaDetect}_{\lambda}$ satisfies
@@
Appendix C Comparison between our surrogate algorithm and INCREASE-c [18]
@@
To evaluate the comparative impact, we replicate the INCREASE-c experimental framework on the same credit card dataset. An isolation forest is trained with a training sample selected uniformly at random from the set of true nulls; a calibration subset of strictly genuine transactions is randomly selected from the null data, after which a test sample is formed by combining the remaining null data with a random subset of fraudulent transactions. The test sample is transformed to $p$-values, and the BH procedure (with a level of $0.1$) is applied to identify the set of fraudulent transactions. (The full details of the experiment can be found in [18].) We execute INCREASE-$c$ against these $p$-values and their ground-truth labels across $100$ simulations for each attack budget $c=1,5,10,20$. The results are reported in Table $6$. The high average value of the selected $p$-values (i.e., $\mathbb{E}[p_{\text{selected}}]$) confirms that INCREASE-c targets data points far from the (average) BH cutoff (i.e., the last column in Table IX). Due to the drastic change in perturbed $p$-values, there is a clear impact on the resulting FDR and total rejection count.
TABLE IX: Experiment E.1: INCREASE-c results: perturbation and cutoff
$c$
original
FDR
INCREASE-c
FDR
original
$\mathbb{E}[R]$
INCREASE-c
$\mathbb{E}[R]$
$\mathbb{E}[p_{\text{selected}}]$
average
BH cutoff
@@
Appendix D Additional Experiments
Synthetic data generation: We generate two types of data: null samples from distribution $P_{0}$ and non-null samples from distribution $P_{1}$. We let $d=20$ for all the synthetic data.
Independent Gaussian: We consider $P_{0}=\mathcal{N}(0,I_{d})$ and $P_{1}=\mathcal{N}(\mu,I_{d})$, where $\mu\in\mathbb{R}^{d}$ is a sparse mean shift vector: the first five coordinates are set to $\sqrt{2\log(d)}$ and the remaining coordinates are zero.
Non-Gaussian: We let the first two coordinates of nulls and non-nulls be drawn independently from Beta distributions: $P_{0}:(X_{1},X_{2})\sim\text{Beta}(5,5)$ and $P_{1}:(X_{1},X_{2})\sim\text{Beta}(1,3)$. The remaining coordinates are drawn i.i.d. from $\text{Beta}(1,1)$ under both $P_{0}$ and $P_{1}$.
@@ Appendix D Additional Experiments
$\Sigma=c\mathbf{1}\mathbf{1}^{\top}+(b^{2}-c)I_{d},\quad\text{ where }\mathbf{1}=[1,\dots,1]^{\top}\in\mathbb{R}^{d}.$
We define the null and non-null distributions as $P_{0}=\mathcal{N}(a\mathbf{1},\Sigma)$ and $P_{1}=\mathcal{N}((a+\delta)\mathbf{1},\Sigma)$, where $\delta>0$ introduces a mean shift across all coordinates. Thus $P_{0}$ and $P_{1}$ share the same exchangeable covariance structure but differ in their mean vectors.
Experiment A.1: Mismatched score function configurations. We investigate both oracle and surrogate attack performance when using different model architectures and parameters for the score functions, with attack size $m_{a}=200$. This experiment comprises distinct configurations for each attack type:
@@
    RF–NN: AdaDetect score function $\tilde{s}(z)$ uses RF, attacker score function g(z){g}(z) uses NN.
@@
    RF–RF: Both score functions use RFs with the same hyperparameters.
TABLE X: Experiment A.2: FDR + RF-NN
@@
TABLE XII: Experiment A.3: Power + NN
@@
Experiment A.2: Real-world data with NN models. This experiment employs NN architectures for both score functions on the four real-world datasets, evaluating both oracle and surrogate attack performance with attack size $m_{a}=200$. This allows us to evaluate how each attack type performs when both the target model and attacker use NN-based approaches on realistic data distributions. We compare boundary attack with our default HSJA attack under both oracle and surrogate attack schemes, using identical RF architectures for both score functions with $m_{a}=200$. The result shows that both HSJA and boundary attack are successful at increasing the FDR for NN models in real-world data.